OpenStudy (anonymous):

given that log a (4)=2.5 and log a(6)=3.23, what is the value of log a(12)?

OpenStudy (anonymous):

@Directrix

OpenStudy (anonymous):

@amistre64

OpenStudy (amistre64):

isnt 12 = 6*2?

OpenStudy (amistre64):

of 12 = 4*3?

OpenStudy (amistre64):

what methods of approach do you know of?

OpenStudy (amistre64):

log(2^2) = 2 log(2) log(6*2) = log(6)+log(2)

OpenStudy (anonymous):

well you have to use log 4 and log 6..

OpenStudy (anonymous):

only those values

OpenStudy (amistre64):

doesnt 4 = 2^2 ?

OpenStudy (amistre64):

log(2*2) = log(2) + log(2) = 2 log(2)

OpenStudy (amistre64):

since 2.5 = 2 log(2) we can solve for log(2) and use it to find log(6*2) = log(6)+log(2)

OpenStudy (anonymous):

wait I'm confused so what would the answer be...